OpenStudy (anonymous):

Find the slope-intercept form of the equation of the line that passes through (-5,3) and is parallel to 12x - 3y =10

OpenStudy (anonymous):

well a line parallel to 12x-3y=10 will have a equation of form 12x-3y=c where c is constant different for different lines parallel to 12x-3y=10 so it passes through -5,3 so put -5 as x and 3 as y in the line 12x-3y=c to get c=-60-9=-69 so the equation of line is 12x-3y=-69 or 12x-3y+69=0

OpenStudy (anonymous):

Ok I got it now. Thank you again! (: